light of wavelength 4000 angstorms fall on a surface of cs calculate the energy of photoelectron emitted if critical wavelength for photoelectric effect in cs is 6600angstroms

Grade:11

1 Answers

Bevkoof Singh
10 years ago

energy released = h(f1-f2){here hf2=threshold frequency}

by calculating we get

f1=7.5*10^13

f2=4.5*1013

so E=7.95*10-21